eText-Home

Rohit Khare

November 4, 1994

Figure ??

Part I

It’s ALIVE!

1 Announcing eText5, Release 0.8

I have no energy left to evangelize this stuff. All I can say is: Figure ?? 1 The App (NeXT, Intel, HP-PA) [Figure ??] ^{2} The Source (All of it) [Figure ??]

3 Demo Documents [Figure ??] ^{4} The old Home Page

1See ftp://etext.caltech.edu/pub/eTextEngine/eText5.App.tar.gz 2See http://www.etext.caltech.edu/pub/eTextEngine/eText5.Source.tar.gz 3See ftp://etext.caltech.edu/pub/eTextEngine/demoETFDs.tar.gz 4See http://www.etext.caltech.edu/oldeTextHome.html

Figure 1: Caption. Figure 2: Caption. Figure 3: Caption.

Figure 4: Caption. Figure 5: Caption.